1. Find all solutions to the system x1 + 2x2 + 3x3 = 8 1 x1 - x2 + 3x3 = 5 =>{801_ ) but ke 2. Is a basis for R3? 13 12
Answer #1

Given system x + 2x2 +323=8 - x2 + 3x2 = 5 the system in matsia Elimination form and solve it by Rewrite Gaussian Applying Ro6 avens 10:10-31 - 11:31 = (10.3)~3x(4-0) to - 7-12 =-5 det 2 slaasto. n. The st] [:] [13] 3* Chesty Independent The number o